OpenStudy (anonymous):

can anyone show me the working to prove whether limt (x-1)/x tht approaches 1 exist or not??

OpenStudy (anonymous):

sure

hero (hero):

first rewrite it as 1 - 1/x

OpenStudy (anonymous):

lim f(x) =(x-1)/x x->1 =0/1 =0

hero (hero):

If you replace x with 1, then 1-1 = 0

hero (hero):

So the limit approaches zero when x = 1

OpenStudy (anonymous):

pls read the ques back Hero..u are explaining the wrong thing..

hero (hero):

No, I'm not

OpenStudy (anonymous):

EvrydayAzn...can show me the working for the left hand n right hand limit??

hero (hero):

Left and right hand? That's not necessary for this

OpenStudy (anonymous):

since its a first degree function, that method is not necessay